From the given table calculate TR, MR and AR and highlight the relationship between total revenue (TR) and marginal revenue (MR). (



1
Expert's answer
2021-08-05T13:07:12-0400
"TR=pQ"

"MR=\frac{\varDelta TR}{\varDelta Q}"

"AR=\frac{\sum TR}{Q}"